OpenStudy (anonymous):

Choose the correct product of (6x - 2)(6x + 2

OpenStudy (anonymous):

36 x^2-4

OpenStudy (anonymous):

36x^2 - 4

OpenStudy (hoblos):

apply the difference of two squares formula which says (a-b)(a+b) = a^2 - b^2 which will give you (6x - 2)(6x + 2) = (6x)^2 - 2^2 = 36x^2 - 4